Exercise 1-16 The summaries of data from the balance sheet, income statement, and retained earnings statement for two corporations, Walco Corporation and Gunther Enterprises, are presented below for 2017. Determine the missing amounts. Assume all changes in stockholders equity are due to changes in retained earnings. Walco Corporation $120,000 62,000 Gunther Enterprises Beginning of year Total assets $144,000 Total liabilities Total stockholders equity 75,000 End of year 178,000 Total assets Total liabilities Total stockholders equity 129,000 51,000 51,100 Changes during year in retained earnings Dividends 5,400 Total revenues 219,000 Total expenses 160,000 82,000 Click if you would like to Show Work for this question: OpenShow Work
Answer #1

Walco Corporation :

1.Stockholders' equity at the beginning of the year:

=Total assets at beginning - Total liabilities at beginning

=120000-62000

=58000

2.Total assets at the end :

=Stockholders equity at the end +Total liabilities at the end

=51100+129000

=180,100

3. Dividends :

(a) Stockholders' equity at the beginning of the year=58000

(b) Net profit =Revenues - Expenses

=219000-160000

=59,000

(c) Stockholders equity at the end =51100

>>Dividends=(a) +(b) - (c)

=58000+59000-51100

=65,900

Gunther Enterprises:

1.Total liabilities at beginning :

=Total assets at beginning - Stockholders' equity at the beginning

=144000 - 75000

=69,000

2.Stockholders equity at the end:

=Total assets at the end - Total liabilities at the end

=178000-51000

=127,000

3.Total revenues :

(a) Stockholders' equity at the beginning of the year=75000

(b)Stockholders' equity at the end =127000

(c) Dividends=5400

(d)Net profit =(c) +(b) - (a)

=5400+127000-75000

=57,400

(e) Total revenue =Net profit +Total expenses

=57400+82000

=139,400

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW.PLEASE RATE MY ANSWER BY HITTING ??. THANK YOU